The Measures of two supplementary angles are in The ratio 5:7 what is The measure of The large side?
Found 2 solutions by stanbon, jpvn2015:
Answer by stanbon(75887)

The Measures of two supplementary angles are in The ratio 5:7 what is The measure of The large side?
-------
5:7 is the same as 5x:7x where 7x is the larger side.
---
Equation:
5x + 7x = 180
12x = 180
x = 15
-----
larger angle = 7x = 15 = 105 degrees
====================================
Cheers,
Stan H.
====================


Answer by jpvn2015(54)

Supplementary angle = 180 degrees
5:7
The large side is 7/12 of 180 = 7x180/12 = 105
Thus, 105 is your answer.